OpenStudy (anonymous):

5 / x - 5 = x / x - 5 - 5/4

OpenStudy (anonymous):

Solve for x and determine if the solution is extraneous or not?

OpenStudy (jiteshmeghwal9):

\[\frac{5}{x-5}=\frac{x}{x-5}\]x-5 will be cancelled & u will gt 'x' :)

OpenStudy (anonymous):

solve them by crosss multiplication

OpenStudy (anonymous):

But there is the minus 5/4 at the end ? It can't cancel out.?

OpenStudy (anonymous):